Use the definition to prove lim x +4 - X -1 2x + 1
Answer #1

on. Use the definition to prove lim - x+4. 2x + Proof by let fel=x+4 0 2x+1 Here, the domain 3 of fail is 1R - 123 and -3 tsi.e. eim set = -3 proved